What is the WAC and WAM of a portfoli containing the four mortgaes described in the following table:

Mortgage Remaining Principal Balance Coupon rare Remining maturity

A 110,000 8.00 340

B 80,000 6.10 355

C 150,000 7.50 325

D 160,000 6.80 330
Answer #1

Following is the table of calculations of weighted average coupon WAC and weighted average maturity WAM:

Mortgage Remaining principal balance Coupon rate Remaining maturity Weights Weighted coupon Weighted maturity
A 1,10,000 8 340 110000/500000 =0.22 8 x 0.22 =1.76 340 x 0.22 =74.8
B 80,000 6.1 355 80000/500000 =0.16 6.1 x 0.16 =0.976 355 x 0.16 =56.8
C 1,50,000 7.5 325 150000/500000 =0.3 7.5 x 0.3 =2.25 325 x 0.3 =97.5
D 1,60,000 6.8 330 160000/500000 =0.32 6.8 x 0.32 =2.176 330 x 0.32 =105.6
Total 5,00,000 0.22+0.16+0.30+0.32 = 1 1.76+0.976+2.25+2.176 = 7 74.8+56.8+97.5+105.6 = 335

So the WAC = 7% and WAM = 335